OpenStudy (anonymous):

~~~~~~~FAN AND MEDAL!~~~~~~~~~~~~~~~~~~~~ Carroll High School has a total of 94 boys and girls who play sports. If the number of boys is 9 more than four times the number of girls, how many boys play sports at this high school? 17 23 77 92 I think its's 77 :/

OpenStudy (anonymous):

94 boys and girls play sports: b + g = 94 The number of boys is 9 more than four times the number of girls. b = 4g + 9 b + g = 94 (4g + 9) + g = 94 5g = 85 g = 85/5 g = 17 b + 17 = 94 b = 77 Check: 4 * 17 + 9 = 68 + 9 = 77 You are correct!

OpenStudy (anonymous):

YAY! thank You Connor!!

OpenStudy (anonymous):

You're welcome!